OpenStudy (elisaneedshelp):

Which of the following is an equation of the line that passes through (3, −5) and has a slope of 2?

OpenStudy (texaschic101):

y = mx + b slope(m) = 2 (3,-5) x = 3 and y = -5 now we sub and solve for b, the y intercept. -5 = 3(2) + b -5 = 6 + b -5 - 6 = b -11 = b your equation is : y = 2x - 11 any questions ? This could have also been solved using the point slope formula.

OpenStudy (elisaneedshelp):

Lol you didnt give me time to post the answer choices

OpenStudy (elisaneedshelp):

A. y + 5 = 2(x − 3) B. y − 5 = 2(x + 3) C. y − 5 = 3(x − 2) D. y + 5 = 3(x − 2)

OpenStudy (texaschic101):

oh.....got to use point slope...here we go y - y1 = m(x - x1) slope(m) = 2 (3,-5) x1 = 3 and y1 = -5 now we sub y - (-5) = 2(x - 3) = y + 5 = 2(x - 3) ==> here you go I was just about to say in the above post that I do not use point slope form unless specifically asked...lol....I guess this one specifically asked...

OpenStudy (texaschic101):

if you have any questions please ask. Sometimes the positive and negative signs can trick you if you are not careful.

OpenStudy (elisaneedshelp):

yeah its just that i hadnt done this in so long but i got what you did there

OpenStudy (texaschic101):

good to hear. It is not really that hard as long as you post the right number in the right spot and pay attention to your signs